OpenStudy (anonymous):

(Algebra II) x/a - 5 = b Solve for X, find any restrictions. Please help!!

OpenStudy (jack1):

how is it? \[\frac{ x }{ a } - 5 = b \rightarrow or \rightarrow \frac{ x }{ a - 5} = b\] ???